OpenStudy (drummergirl7):

Help ill medal and fan! :) A bowl contains 15 granola bars and 10 fruit bars. What is the ratio of granola bars to fruit bars? Enter your answer in simplest fractional form in the box.

OpenStudy (drummergirl7):

@Eiwoh @steve816 @jabez177

OpenStudy (drummergirl7):

hey!

OpenStudy (steve816):

Heyyyy

OpenStudy (steve816):

The ratio would be 15 to 10 Or you can write it like 15:10

OpenStudy (drummergirl7):

How r u guys!! :)

OpenStudy (steve816):

Good, thank you :) What about you?

OpenStudy (blazeryder):

\[\frac{ 15 }{ 10 }\]=\[1\frac{ 5 }{ 10 }\]This can be reduced to \[1\frac{ 1 }{ 2 }\]

OpenStudy (drummergirl7):

good! just hangin out XD

OpenStudy (steve816):

You are write, you can reduce it to \[\Huge 3:2\]

OpenStudy (steve816):

*right

OpenStudy (drummergirl7):

3 :2 is that my answer :)

OpenStudy (steve816):

Yup!

OpenStudy (drummergirl7):

yay!